# Mask-Enhanced Segment Anything Model for Tumor Lesion Semantic Segmentation

### v1

@@ Abstract
Tumor lesion segmentation on CT or MRI images plays a critical role in cancer diagnosis and treatment planning. Considering the inherent differences in tumor lesion segmentation data across various medical imaging modalities and equipment, integrating medical knowledge into the Segment Anything Model (SAM) presents promising capability due to its versatility and generalization potential. Recent studies have attempted to enhance SAM with medical expertise by pre-training on large-scale medical segmentation datasets. However, challenges still exist in 3D tumor lesion segmentation owing to tumor complexity and the imbalance in foreground and background regions. Therefore, we introduce Mask-Enhanced SAM (M-SAM), an innovative architecture tailored for 3D tumor lesion segmentation. We propose a novel Mask-Enhanced Adapter (MEA) within M-SAM that enriches the semantic information of medical images with positional data from coarse segmentation masks, facilitating the generation of more precise segmentation masks. Furthermore, an iterative refinement scheme is implemented in M-SAM to refine the segmentation masks progressively, leading to improved performance. Extensive experiments on seven tumor lesion segmentation datasets indicate that our M-SAM not only achieves high segmentation accuracy but also exhibits robust generalization.
@@ 2.1 Overall Review
The overall architecture of our M-SAM is illustrated in Fig. 1. M-SAM is built upon the architecture of SAM-Med3D [23], which has been pre-trained on large-scale medical image segmentation datasets to obtain general medical knowledge. It is composed of a ViT-based 3D image encoder to extract 3D image embeddings, a 3D prompt encoder to extract prompt embeddings, and a lightweight 3D mask decoder to predict segmentation results. As shown in Fig. 1, given a 3D image (CT or MRI) $I\in\mathbb{R}^{C\times H\times W\times Dp}$, where $C$ represents the number of channels, $H$ the height, $W$ the width, and $Dp$ the depth of the image, determined by the number of modalities of the image (e.g., the T1 and T2 modalities of MRI) and a randomly-initialized point as the initial point prompt. We first process them with SAM image encoder and SAM prompt encoder respectively, to obtain the initial image embedding $E_{I}^{0}\in\mathbb{R}^{N_{I}\times D}$ and point embedding $E_{P}^{0}\in\mathbb{R}^{N_{P}\times D}$. Here, $N_{I}$ and $N_{P}$ denote the dimensions of the image and point embeddings, respectively, and $D$ represents the feature dimensionality. We also initialize a null mask $M\in\mathbb{R}^{C\times H\times W\times Dp}$ consisting of all zeros and feed it into a mask encoder to obtain the initial mask embedding $E_{M}^{0}\in\mathbb{R}^{N_{I}\times D}$. Afterward, the image embedding $E_{I}^{0}$ and mask embedding $E_{M}^{0}$ are fed into our Mask-Enhanced Adapter to obtain the updated image embedding $\hat{E}_{I}^{0}$. By feeding $\hat{E}_{I}^{0}$ and $E_{P}^{0}$ into the SAM mask decoder, coarse segmentation mask $M^{1}$ can be produced, which can also be used to guide the update of image embeddings in MEA for the next stage of segmentation refinement. The image embeddings are also updated in the iterative process accordingly. Through multiple stages of iterative refinement, it is possible to improve the segmentation masks in a coarse-to-fine manner continuously, thus boosting segmentation performance.
@@ 3.1 Experimental Setting
Implementation Details. We use the AdamW optimizer with an initial learning rate of $48e-$ 4 and train for $200$ epochs and the batch size was set to $4$, with a weight decay of $0.1$. In our dataset transform process, we employ a crop-or-pad strategy to standardize all images to a resolution of $128\times 128\times 128$. This involves zero-padding for images with any dimension falling short of $128$ and applying trilinear interpolation to resize images that exceed the specified dimensions. As for the data augmentation, we use RandomFlip along all three spatial axes and perform ZNormalization on each medical image data. All the experiments are implemented in PyTorch and trained on one NVIDIA Tesla V100 GPU.
@@ 3.2 Main Result
Comparison with State-of-the-Art methods. As shown in the Table 1, our method outperforms all other methods on the aforementioned datasets. Notably, compared to the existing methods, our method achieves an average improvement of approximately $2$%. This demonstrates the effectiveness of our method across multiple types of tumor lesion datasets. Additionally, it can be observed from the table that U-Net-based methods tend to perform better on individual datasets. However, the performance of such methods varies significantly across different datasets. For example, SwinUNETR [10] achieves a Dice score of over $90$% on BraTS21, but only achieves a score of $70.7$% on the Pancreas dataset, which is nearly $10$% lower than our method. For other SAM-based methods, the Dice scores are generally lower than those of UNet-based methods. In terms of the number of parameters, our model has fewer parameters than the average of the methods shown in the table, and the tunable parameter count also ranks fourth among these methods. These experimental results indicate that our method can surpass these state-of-the-art methods with lower computational costs, effectively leveraging the advantages of SAM in the field of tumor lesion segmentation.

### v2

@@ Abstract
Tumor lesion segmentation on CT or MRI images plays a critical role in cancer diagnosis and treatment planning. Considering the inherent differences in tumor lesion segmentation data across various medical imaging modalities and equipment, integrating medical knowledge into the Segment Anything Model (SAM) presents promising capability due to its versatility and generalization potential. Recent studies have attempted to enhance SAM with medical expertise by pre-training on large-scale medical segmentation datasets. However, challenges still exist in 3D tumor lesion segmentation owing to tumor complexity and the imbalance in foreground and background regions. Therefore, we introduce Mask-Enhanced SAM (M-SAM), an innovative architecture tailored for 3D tumor lesion segmentation. We propose a novel Mask-Enhanced Adapter (MEA) within M-SAM that enriches the semantic information of medical images with positional data from coarse segmentation masks, facilitating the generation of more precise segmentation masks. Furthermore, an iterative refinement scheme is implemented in M-SAM to refine the segmentation masks progressively, leading to improved performance. Extensive experiments on seven tumor lesion segmentation datasets indicate that our M-SAM not only achieves high segmentation accuracy but also exhibits robust generalization. The code is available at https://github.com/nanase1025/M-SAM.
@@ 2.1 Overall Review
The overall architecture of our M-SAM is illustrated in Fig. 1. M-SAM is built upon the architecture of SAM-Med3D [22], which has been pre-trained on large-scale medical image segmentation datasets to obtain general medical knowledge. It is composed of a ViT-based 3D image encoder to extract 3D image embeddings, a 3D prompt encoder to extract prompt embeddings, and a lightweight 3D mask decoder to predict segmentation results. As shown in Fig. 1, the inputs are a 3D image (CT or MRI) $I\in\mathbb{R}^{C\times H\times W\times Dp}$, where $C$ represents the number of channels, $H$ the height, $W$ the width, and $Dp$ the depth (determined by the number of image modalities, e.g., the T1 and T2 modalities of MRI), and a randomly-initialized point as the initial point prompt. We first process them with SAM image encoder and SAM prompt encoder respectively, to obtain the initial image embedding $E_{I}^{0}\in\mathbb{R}^{N_{I}\times D}$ and point embedding $E_{P}^{0}\in\mathbb{R}^{N_{P}\times D}$. Here, $N_{I}$ and $N_{P}$ denote the dimensions of the image and point embeddings, respectively, and $D$ represents the feature dimensionality. We also initialize a null mask $M\in\mathbb{R}^{C\times H\times W\times Dp}$ consisting of all zeros and feed it into a mask encoder to obtain the initial mask embedding $E_{M}^{0}\in\mathbb{R}^{N_{I}\times D}$. Afterward, the image embedding $E_{I}^{0}$ and mask embedding $E_{M}^{0}$ are fed into our Mask-Enhanced Adapter to obtain the updated image embedding $\hat{E}_{I}^{0}$. By feeding $\hat{E}_{I}^{0}$ and $E_{P}^{0}$ into the SAM mask decoder, coarse segmentation mask $M^{1}$ can be produced, which can also be used to guide the update of image embeddings in MEA for the next stage of segmentation refinement. The image embeddings are also updated in the iterative process accordingly. Through multiple stages of iterative refinement, it is possible to improve the segmentation masks in a coarse-to-fine manner continuously, thus boosting segmentation performance.
@@ 2.3 Iterative Refinement
Following SAM-Med3D [22], to simulate the clinical scenario of interactive segmentation, one point per iteration is randomly sampled: from the foreground in the 1st iteration, and from the error region between the coarse mask and ground truth in the subsequent 9 iteration, totaling 10 points in $10$ iterations. The updated prompt embedding $E_{P}$ is thus generated by feeding the newly sampled point into the prompt encoder at each iteration. This point sampling strategy is used for both training and inference in our experiments, while in real clinical use, it operates interactively with physicians.
@@ 3.1 Experimental Setting
Implementation Details. We use the AdamW optimizer with an initial learning rate of $48e-$ 4 and train for $200$ epochs and the batch size was set to $4$, with a weight decay of $0.1$. In our dataset transform process, we employ a crop-or-pad strategy to standardize all images to a resolution of $128\times 128\times 128$. This involves zero-padding for images with any dimension falling short of $128$ and applying cropping for dimensions exceeding $128$. As for the data augmentation, we use RandomFlip along all three spatial axes and perform z-score normalization on each medical image data. All the experiments are implemented in PyTorch and trained on one NVIDIA Tesla V100 GPU.
@@ 3.2 Main Result
Comparison with State-of-the-Art methods. As shown in the Table 1, our method outperforms all other methods on the aforementioned datasets. Notably, compared to the existing methods, our method achieves an average improvement of approximately $2$%. This demonstrates the effectiveness of our method across multiple types of tumor lesion datasets. Additionally, it can be observed from the table that U-Net-based methods tend to perform better on individual datasets. However, the performance of such methods varies significantly across different datasets. For example, SwinUNETR [10] achieves a Dice score of over $90$% on BraTS21, but only achieves a score of $70.7$% on the Pancreas dataset, which is nearly $10$% lower than our method. For other SAM-based methods, the Dice scores are generally lower than those of UNet-based methods. In terms of the number of parameters, our model has fewer parameters than the average of the methods shown in the table, and the tunable parameter count also ranks fourth among these methods. These experimental results indicate that our method can surpass these state-of-the-art methods with lower computational costs, effectively leveraging the advantages of SAM in the field of tumor lesion segmentation. Our results in Table 1 are averaged across 5 runs, with Dice STD less than 0.3% on all datasets.
@@ 4 Conclusion
Acknowledgments. This research is supported in part by National Natural Science Foundation of China (NO. 62122010, U23B2010 and NO. 62322608), Zhejiang Provincial Natural Science Foundation of China (Grant NO. LDT23F02022F02), Key Research and Development Program of Zhejiang Province (Grant 2022C01082) and Beijing Natural Science Foundation (NO. L231011).